OpenStudy (reganmckay32):

A dolldoll sold for ​$277 in 1980 and was sold again in 1986 for $434. Assume that the growth in the value V of the​ collector's item was exponential. ​a) Find the value k of the exponential growth rate. V subscript o=277.

OpenStudy (518nad):

hi okay, they want u to assume there was some exponential growth rate per year*

OpenStudy (518nad):

i am not sure if you've been introduced to writing exponential formulas with the base e in the form Y=Ae^(kt)

OpenStudy (518nad):

if you are not, you can still use the same logic as before, suppose there is some growth rate per year

OpenStudy (518nad):

if u have a growth rate g for 6 years we have 277(1+g)^6=434 we can solve for g

OpenStudy (518nad):

or with this equation Y=Ae^(kt) 434=277e^(k*6) solve for k

OpenStudy (518nad):

this k is going to be different from the g btw, because its a continuous growth rate vs incremental growth rate

OpenStudy (reganmckay32):

We do write with base e! :)

OpenStudy (518nad):

okay solve for k there

OpenStudy (518nad):

u can use this Y=xe^(kt) then (Y/x) = e^(kt) ln(Y/x)=kt k=(ln(Y/x))/t

OpenStudy (mathmale):

Using "x" as a constant here is not wise. I propose you use the exponential growth formula\[A=A _{0}Ce ^{kt}\]

OpenStudy (mathmale):

where A would represent the value of the doll in 1986, A_0 is the value of the doll in 1980, and k is the exponential growth constant. (Cross out the "C" in the above equation; do not use it.)

OpenStudy (mathmale):

t represents the number of years that have elapsed. What would the value of t be here?